OpenStudy (sstradic):

MEDAL!!! A glass vase is shaped like a cylinder and has a radius of 2.5 inches. The vase holds 170 cubic inches of water. How tall is the vase? Use 3.14 for pi. Round your answer to the nearest hundredth. 8.66 inches 9.78 inches 10.35 inches 12.50 inches

Vocaloid (vocaloid):

vollume = pi * radius^2 * height you are given the radius and volume. use the equation to solve for height

Vocaloid (vocaloid):

you are allowed to ask questions if you do not understand

OpenStudy (sstradic):

WOULD THE VOLUME BE 170?

Vocaloid (vocaloid):

yes! do you know how to solve for height now?

OpenStudy (sstradic):

no :(

Vocaloid (vocaloid):

let's take a look at our equation volume = pi * radius^2 * height now let's replace "volume" with 170 and "radius" with 2.5

Vocaloid (vocaloid):

170 = pi * (2.5)^2 * height

OpenStudy (mathmale):

Yes, the volume of the cylinder is 170 cubic inches. Your job is to solve for 'height:' 170 = pi * (2.5)^2 * height

OpenStudy (mathmale):

170 cubic inches = pi * (6.25 square inches) * height. Solve for height.

Vocaloid (vocaloid):

@sstradic since we have "height" multiplied by pi and 6.25, we want to "undo" the multiplication to get "height" by itself what's the opposite of multiplication?

OpenStudy (sstradic):

division(:

Vocaloid (vocaloid):

good, so we divide both sides by 6.25 and pi

Vocaloid (vocaloid):

therefore height = 170/(6.25 * pi) = ?

OpenStudy (sstradic):

I got 8.66 thank you so much!